# Born in the Dark: The Catastrophic Collapse of Fuzzy Dark Matter Solitons as the Origin of Little Red Dots

### v1

@@ Abstract
JWST surveys have uncovered a population of compact, red sources (“Little Red Dots,” LRDs) at $z\gtrsim 5$ that exhibit broad Balmer emission yet remain X-ray faint, implying heavy obscuration with $N_{\rm H}\gtrsim 10^{24}\,{\rm cm^{-2}}$. We propose that LRDs may trace a short-lived, obscured phase associated with rapid baryonic inflow inside the deep solitonic cores of fuzzy dark matter (FDM) halos. Combining the soliton size scaling with (i) the observed compact radii ($r_{e}\sim 30$100100 pc) and (ii) the requirement that Compton-thick columns be achievable within a region of order the core radius, we find that particle masses $m\sim{\rm few}\times 10^{-22}\,{\rm eV}$ are plausible for soliton masses $M_{s}\sim 10^{8}$$10^{9}\,M_{\odot}$; we adopt $m_{22}=2$ as a fiducial choice. A conservative mass-budget estimate for the obscuring column, together with isothermal hydrostatic stratification, indicates that configurations reaching $N_{\rm H}\gtrsim 10^{24}$$10^{25}\,{\rm cm^{-2}}$ require densities for which radiative losses (cooling and/or diffusion) occur faster than the dynamical time, suggesting that a long-lived static hot atmosphere is unlikely (an “Opacity Crisis”) and that rapid inflow or radiation-pressure-driven evolution is favored. Using $512^{3}$ pseudo-spectral Schrödinger–Poisson simulations of idealized soliton mergers, we illustrate that compact, high-density soliton cores can form via violent relaxation under representative scalings. We discuss observational implications and tests, and outline the need for future radiation-hydrodynamic modeling to predict demographics and detailed spectra.
@@ I Introduction
The origin of supermassive black holes (SMBHs) residing in the centers of galaxies remains one of the most puzzling questions in modern astrophysics. The detection of quasars at redshift $z>7$ with black hole masses exceeding $10^{9}M_{\odot}$ (Mortlock et al., 2011; Banados et al., 2018) places stringent timing constraints on their growth. Standard scenarios, such as light seeds from Pop III stars or heavy seeds from Direct Collapse Black Holes (DCBH), face significant challenges regarding accretion rates and rare environmental requirements (Bromm, 2013).
@@
Fuzzy Dark Matter (FDM) provides a novel framework for addressing these issues. FDM halos host a stable, dense ground-state solution known as a soliton (Schive et al., 2014a). In this paper, we propose that LRDs are the direct observational signatures of the catastrophic breakdown of the hydrostatic atmosphere triggered by FDM soliton thermodynamics. We identify a regime of “Opacity Crisis” where the static atmosphere becomes unphysical. We present analytic derivations and 3D numerical simulations to show that: (1) deep solitonic potential wells form robustly, and (2) gas within these wells must undergo rapid cooling and inflow.
@@ II.1 Soliton Scaling and Parameter Constraints
where $m_{22}=m/10^{-22}\text{ eV}$. We constrain $m_{22}$ by requiring consistency with LRD observables: compact sizes ($r_{\rm e}\sim 30$100100 pc) and Compton-thick obscuration ($N_{\rm H}>10^{24}\rm{cm}^{-2}$). For a typical LRD host halo, the core-halo relation implies $M_{\rm s}\sim 10^{9}M_{\odot}$.
As shown in Figure 1, we identify the parameter space where FDM is consistent with observations. To ensure robustness, we also perform a sensitivity analysis regarding the baryon loading parameters. We note that the derived constraints depend on the assumption that the observed effective radius $r_{e}$ traces the soliton core scale $r_{c}$ (i.e., $\eta_{r}\equiv r_{e}/r_{c}\sim\mathcal{O}(1)$). If $r_{e}$ corresponds to a larger scattering photosphere, the constraints would shift; however, for our fiducial assumptions, a window around $m_{22}\approx 2$ emerges.
Figure 1: Constraints on boson mass ($m_{22}$). The viable region is defined by the intersection of the observed LRD size range (green band, $r_{e}\sim 30$100100 pc) and the requirement that Compton-thick columns be achievable within a compact core. The red hatched region marks parameters for which even a maximally loaded core ($f_{g}=1,\xi=1$) fails to reach $N_{\rm H}=10^{24}\,{\rm cm^{-2}}$. The opacity constraint curves assume a soliton mass scale of $M_{s}\sim 10^{9}M_{\odot}$ (typical of LRD hosts). We treat the observed $r_{e}$ as a proxy for the soliton core radius $r_{c}$; if $r_{e}$ traces a larger photosphere, the constraints would shift. To address uncertainties in baryon physics, the orange shaded area shows the sensitivity range: under conservative assumptions ($f_{g}=0.1,\xi=3$), the exclusion region expands. Under these working assumptions, a parameter window around $m_{22}\sim\text{few}$ is favored.
@@ II.2 Derivation of the Opacity Crisis
We ask whether a quasi-static baryonic atmosphere inside a soliton can simultaneously (i) produce Compton-thick obscuration and (ii) remain pressure-supported on a dynamical time. Rather than assuming a unique static solution, we explicitly parameterize the baryon loading and show that the regime required for obscuration generically implies rapid radiative losses, invalidating the long-lived hydrostatic picture.
@@
Let a gas mass $M_{g}=f_{g}f_{b}M_{s}$ be loaded into an obscuring region of radius $r_{\rm out}=\xi r_{c}$, where $f_{g}\leq 1$ is the loaded fraction of the cosmic baryon budget, $f_{b}\simeq 0.16$, and $\xi\sim 1$–few. A geometry-limited characteristic column is
$N_{\rm H}\simeq\frac{M_{g}}{\pi r_{\rm out}^{2}\,\mu_{\rm H}m_{p}}=\frac{f_{g}f_{b}M_{s}}{\pi(\xi r_{c})^{2}\,\mu_{\rm H}m_{p}},$ (2)
with $\mu_{\rm H}\simeq 1.4$ accounting for He. Using the soliton scaling (Eq. 1), this implies that for fixed $(f_{g},\xi)$ the achievable column increases rapidly with deeper, more compact cores. In Fig. 1 we define the “Too Transparent” region by $N_{\rm H,max}(f_{g}=1,\xi=1)<10^{24}\,{\rm cm^{-2}}$. As shown by the orange band in Fig. 1, adopting conservative values ($f_{g}=0.1,\xi=3$) shifts this threshold, but $m_{22}\approx 2$ remains a plausible solution for $M_{s}\sim 10^{9}M_{\odot}$. We note that Eq. (2) is a conservative geometric estimate; gas clumping would increase the effective line-of-sight column, while feedback might reduce $f_{g}$.
@@
which can be much shorter than the dynamical (free-fall) time $t_{\rm ff}\sim(G\bar{\rho})^{-1/2}$.
@@
In an optically thick core, radiative losses may be limited by diffusion rather than by optically thin cooling. A conservative lower bound is the diffusion time due to electron scattering,
$\begin{split}t_{\rm diff,es}\sim&\frac{\tau_{T}r_{\rm out}}{c}=\frac{\sigma_{T}N_{H}r_{\rm out}}{c}\\ \simeq&1.6\times 10^{3}\,{\rm yr}\,\left(\frac{\tau_{T}}{10}\right)\left(\frac{r_{\rm out}}{50\,{\rm pc}}\right),\end{split}$ (5)
which is typically shorter than $t_{\rm ff}$ in the parameter range where Compton-thick columns are attainable.
If the gas cools and becomes dusty, UV/optical photons can be efficiently reprocessed into the IR, and the relevant mass opacity can be much larger than the electron-scattering opacity $\kappa_{\rm es}\simeq 0.34~{\rm cm^{2}\,g^{-1}}$. In that regime, an upper-range diffusion estimate is
$t_{\rm diff,IR}\sim\frac{\tau_{\rm IR}r_{\rm out}}{c},\quad\tau_{\rm IR}=\kappa_{\rm IR}\Sigma,\quad\Sigma=\mu_{H}m_{p}N_{H},$ (6)
@@
$\begin{split}t_{\rm diff,IR}\simeq&3.8\times 10^{4}\,{\rm yr}\,\left(\frac{\kappa_{\rm IR}}{10\,{\rm cm^{2}\,g^{-1}}}\right)\\ &\times\left(\frac{N_{H}}{10^{25}\,{\rm cm^{-2}}}\right)\left(\frac{r_{\rm out}}{50\,{\rm pc}}\right).\end{split}$ (7)
@@
We use “Opacity Crisis” to denote the breakdown of a long-lived, hot hydrostatic atmosphere inside the soliton: in the same parameter space where $N_{H}\gtrsim 10^{24}\,{\rm cm^{-2}}$ is achievable within $r_{\rm out}\sim O(r_{c})$, the effective radiative-loss or energy-redistribution time $t_{\rm rad}\equiv\max(t_{\rm cool},t_{\rm diff,es})$ is not $\gg t_{\rm ff}$. The system therefore cannot remain as a quasi-static hot atmosphere; it must evolve dynamically, either via rapid cooling-driven collapse ($t_{\rm cool}\ll t_{\rm ff}$) or into a radiation-pressure-supported optically thick inflow/envelope if IR trapping becomes important ($t_{\rm diff,IR}\gtrsim t_{\rm ff}$). In either case, the initial static HSE picture is invalid in the regime relevant for LRD-like obscuration.
@@ II.3 Timescale analysis and inflow rate
Figure 2 summarizes the timescale hierarchy for our fiducial choice $m_{22}=2.0$. In the parameter space where Eq. (2) allows $N_{\rm H}\gtrsim 10^{24}\,{\rm cm^{-2}}$, radiative losses are rapid compared to dynamical support.
A direct implication is that the effective sound speed can drop substantially during collapse, enhancing the capture rate. For a black hole of mass $M_{\rm BH}$, the Bondi rate is
@@
The ratio to the Eddington rate is
@@
indicating that inflow at the capture scale can exceed $\dot{M}_{\rm Edd}$ in dense cores. Granule-driven fluctuations (Quantum Bondi Boost) may further modulate this (Chiu et al., 2025).
Figure 2: The instability region. Comparison of free-fall ($t_{\rm ff}$) and cooling ($t_{\rm cool}$) timescales versus soliton mass for $m_{22}=2.0$. The shaded region indicates $t_{\rm cool}<t_{\rm ff}$, where a long-lived, optically-thin hot hydrostatic atmosphere cannot persist. The system is driven either toward rapid inflow/collapse (if radiative losses escape efficiently) or toward a radiation-pressure dominated, optically thick configuration (if radiation is trapped).

### v2

@@ Abstract
JWST surveys have uncovered a population of compact, red sources (“Little Red Dots,” LRDs) at $z\geq 5$ that exhibit broad Balmer emission yet remain X-ray faint, implying heavy obscuration with $N_{\rm H}\geq 10^{24}\,{\rm cm^{-2}}$. We propose that LRDs trace a short-lived, obscured phase associated with rapid baryonic inflow inside the deep solitonic cores of fuzzy dark matter (FDM) halos. Combining the soliton size scaling with the observed compact radii ($r_{e}\sim 30$100100 pc), we find that while a particle mass of $m\sim 2\times 10^{-22}$ eV provides a direct match for mature systems, the observed size-mass relation is fully consistent with heavier bosons ($m_{22}\gtrsim 20$, satisfying Lyman-$\alpha$ constraints) if LRDs represent a progenitor phase where the central black hole is still growing ($M_{\rm BH}<M_{c}$). We adopt $m_{22}=2$ as a fiducial baseline to demonstrate the thermodynamic instability. A conservative mass-budget estimate indicates that configurations reaching $N_{\rm H}\geq 10^{24}$$10^{25}\,{\rm cm^{-2}}$ require densities for which radiative losses (cooling and/or diffusion) occur faster than the dynamical time, suggesting that a long-lived static hot atmosphere is unlikely (an “Opacity Crisis”) and that rapid inflow or radiation-pressure-driven evolution is favored. Using $512^{3}$ pseudo-spectral Schrödinger–Poisson simulations of idealized soliton mergers, we illustrate that compact, high-density soliton cores form robustly under representative scalings. We discuss observational implications and tests, and outline the need for future radiation-hydrodynamic modeling to predict demographics and detailed spectra.
@@ II.1 Soliton Scaling and Parameter Constraints
where $m_{22}=m/10^{-22}\text{ eV}$. We constrain $m_{22}$ by requiring consistency with LRD observables: compact sizes ($r_{\rm e}\sim 30$100100 pc) and Compton-thick obscuration ($N_{\rm H}>10^{24}\rm{cm}^{-2}$). For a typical LRD host halo, the core–halo relation implies $M_{\rm s}\sim 10^{9}\,M_{\odot}$ (e.g., Schive et al., 2014b; Chan et al., 2018).
As shown in Figure 1, we identify the parameter space where FDM is consistent with observations. To ensure robustness, we also perform a sensitivity analysis regarding the baryon loading parameters. We note that the derived constraints depend on the assumption that the observed effective radius $r_{e}$ traces the soliton core scale $r_{c}$ (i.e., $\eta_{r}\equiv r_{e}/r_{c}\sim\mathcal{O}(1)$). For the soliton scaling in Eq. (1), one finds $m_{22}\propto r_{c}^{-1/2}\propto\eta_{r}^{1/2}$ at fixed $(r_{e},M_{s})$; thus an uncertainty $\eta_{r}\sim 0.5$33 shifts the inferred $m_{22}$ by a factor of $\sim 0.7$$1.7$. If $r_{e}$ corresponds to a larger scattering photosphere, or if the black hole is in an early progenitor phase where it does not yet trace the full soliton mass ($M_{\rm BH}<M_{c}$), the constraints would shift significantly. However, under the direct mapping assumption ($M_{\rm BH}\approx M_{c}$), a fiducial window around $m_{22}\approx 2$ emerges.
Figure 1: Constraints on boson mass ($m_{22}$) from combining the soliton size scaling (Eq. 1) with an opacity (column) requirement. The vertical axis shows a characteristic physical scale; we compare the observed effective radius range of LRDs (green band, $r_{e}\sim 30$100100 pc) with the soliton core radius predicted for $M_{s}=10^{8}$$10^{9}\,M_{\odot}$ (blue band; dashed boundaries follow $r_{c}\propto m_{22}^{-2}$). The red hatched region (“Too transparent”/“Generic transparency”) marks parameters for which even a maximally loaded core ($f_{g}=1$) confined within $r_{\rm e}=\xi r_{c}$ with $\xi=1$ cannot reach $N_{\rm H}=10^{24}\,{\rm cm^{-2}}$. The opacity boundary is evaluated for a representative $M_{s}\sim 10^{9}\,M_{\odot}$ (typical of LRD hosts). The orange shaded area illustrates sensitivity to baryon-loading uncertainties by adopting conservative values ($f_{g}=0.1$, $\xi=3$). The viable region is the overlap between the observed size band and the non-transparent (Compton-thick-capable) region; The region outside the red-shaded area corresponds entirely to the non-transparent (Compton-thick capable) regime. The “Viable Region” is specifically the parallelogram defined by the overlap between the observed size band and this non-transparent regime. under the direct-mapping assumption ($r_{e}\sim r_{c}$) this favors $m_{22}\sim{\rm few}$, and we adopt $m_{22}=2$ as a fiducial baseline for subsequent figures.
@@ II.2 Derivation of the Opacity Crisis
We ask whether a quasi-static baryonic atmosphere inside a soliton can simultaneously (i) produce Compton-thick obscuration and (ii) remain pressure-supported on a dynamical time. Rather than assuming a unique static solution, we explicitly parameterize the baryon loading and show that the regime required for obscuration generically implies rapid radiative losses, invalidating the long-lived hydrostatic picture. We define the relevant dynamical timescale at radius $r$ as determined by the enclosed mass potential,
$t_{\rm dyn}(r)\equiv\sqrt{\frac{r^{3}}{G\,M_{\rm enc}(r)}},$ (2)
where $M_{\rm enc}(r)$ is the enclosed gravitating mass. In the BH-dominated (“naked”) limit at $r\sim r_{\rm e}$, $M_{\rm enc}\approx M_{\rm BH}$, whereas in the soliton-dominated case $M_{\rm enc}\approx M_{\rm BH}+M_{\rm sol}(<r)\geq M_{\rm BH}$. At fixed physical radius $r\sim r_{\rm e}\sim{\cal O}(r_{c})$ this implies $t_{\rm dyn,sol}\leq t_{\rm dyn,BH}$. Since the instability is diagnosed by radiative losses/redistribution not being slow compared to dynamical support (i.e., $t_{\rm rad}\lesssim t_{\rm dyn}$), the conclusion that a long-lived hydrostatic atmosphere is untenable in the soliton case implies the same conclusion a fortiori in the BH-only limit, where the dynamical time is longer. The role of the soliton is therefore not to “save” a static atmosphere, but to define the characteristic LRD scale and to motivate considering a quasi-spherical, volume-filling configuration at $r\sim 50$100100 pc in the first place. In what follows, we identify the characteristic physical extent of the obscurer/atmosphere directly with the observed effective radius $r_{\rm e}$, and assume $r_{\rm e}\sim{\cal O}(r_{c})$.
@@
Let a gas mass $M_{g}=f_{g}f_{b}M_{s}$ be loaded into an obscuring region of radius $r_{\rm e}=\xi r_{c}$, where $f_{g}\leq 1$ is the loaded fraction of the cosmic baryon budget (we explore $f_{g}\sim 0.1$11 to bracket feedback-driven depletion; e.g., Hopkins et al. 2014), $f_{b}\simeq 0.16$ (Planck Collaboration et al., 2020), and $\xi\sim 1$–few. A geometry-limited characteristic column is
$N_{\rm H}\simeq\frac{M_{g}}{\pi r_{\rm e}^{2}\,\mu_{\rm H}m_{p}}=\frac{f_{g}f_{b}M_{s}}{\pi(\xi r_{c})^{2}\,\mu_{\rm H}m_{p}},$ (3)
with $\mu_{\rm H}\simeq 1.4$ accounting for He. Using the soliton scaling (Eq. (1)), this implies that for fixed $(f_{g},\xi)$ the achievable column increases rapidly with deeper, more compact cores. In Fig. 1 we define the “Generic Transparency” region by $N_{\rm H,max}(f_{g}=1,\xi=1)<10^{24}\,{\rm cm^{-2}}$. As shown by the orange band in Fig. 1, adopting conservative values ($f_{g}=0.1,\xi=3$) shifts this threshold, but $m_{22}\approx 2$ remains a plausible solution for $M_{s}\sim 10^{9}M_{\odot}$. We note that Eq. (3) is a conservative geometric estimate; gas clumping would increase the effective line-of-sight column, while feedback might reduce $f_{g}$.
@@
where $\Lambda(T,Z)$ is the radiative cooling function (we use the tabulations of Sutherland & Dopita 1993), which can be much shorter than the potential-defined dynamical time $t_{\rm dyn}(r_{\rm e})$ (Eq. 2). For reference, a self-gravitating free-fall time would scale as $\sim(G\bar{\rho})^{-1/2}$; in the soliton case the relevant support time is set by the enclosed (BH+soliton) potential.
@@
In an optically thick core, radiative losses may be limited by diffusion rather than by optically thin cooling. A conservative lower bound is the diffusion time due to electron scattering, where $\tau_{T}\equiv\sigma_{T}N_{H}$ is the Thomson optical depth,
$\begin{split}t_{\rm diff,es}\sim&\frac{\tau_{T}r_{\rm e}}{c}=\frac{\sigma_{T}N_{H}r_{\rm e}}{c}\\ \simeq&1.6\times 10^{3}\,{\rm yr}\,\left(\frac{\tau_{T}}{10}\right)\left(\frac{r_{\rm e}}{50\,{\rm pc}}\right),\end{split}$ (6)
which is typically shorter than $t_{\rm dyn}$ in the parameter range where Compton-thick columns are attainable. If the gas cools and becomes dusty, UV/optical photons can be efficiently reprocessed into the IR, and the relevant mass opacity can be much larger than the electron-scattering opacity $\kappa_{\rm es}\simeq 0.34~{\rm cm^{2}\,g^{-1}}$. In that regime, an upper-range diffusion estimate is
$t_{\rm diff,IR}\sim\frac{\tau_{\rm IR}r_{\rm e}}{c},\quad\tau_{\rm IR}=\kappa_{\rm IR}\Sigma,\quad\Sigma=\mu_{H}m_{p}N_{H},$ (7)
@@
$\begin{split}t_{\rm diff,IR}\simeq&3.8\times 10^{4}\,{\rm yr}\,\left(\frac{\kappa_{\rm IR}}{10\,{\rm cm^{2}\,g^{-1}}}\right)\\ &\times\left(\frac{N_{H}}{10^{25}\,{\rm cm^{-2}}}\right)\left(\frac{r_{\rm e}}{50\,{\rm pc}}\right).\end{split}$ (8)
@@
We use “Opacity Crisis” to denote the breakdown of a long-lived, hot hydrostatic atmosphere inside the soliton: in the same parameter space where $N_{H}\gtrsim 10^{24}\,{\rm cm^{-2}}$ is achievable within $r_{\rm e}\sim O(r_{c})$, the effective radiative-loss or energy-redistribution time $t_{\rm rad}\equiv\max(t_{\rm cool},t_{\rm diff,es})$ is not $\gg t_{\rm dyn}$. The system therefore cannot remain as a quasi-static hot atmosphere; it must evolve dynamically, either via rapid cooling-driven collapse ($t_{\rm cool}\ll t_{\rm dyn}$) or into a radiation-pressure-supported optically thick inflow/envelope if IR trapping becomes important ($t_{\rm diff,IR}\gtrsim t_{\rm dyn}$). In either case, the initial static HSE picture is invalid in the regime relevant for LRD-like obscuration.
@@ II.3 Timescale analysis and inflow rate
Figure 2 summarizes the timescale hierarchy for our fiducial choice $m_{22}=2.0$. In the parameter space where Eq. (3) allows $N_{\rm H}\gtrsim 10^{24}\,{\rm cm^{-2}}$, radiative losses are rapid compared to dynamical support. A direct implication is that the effective sound speed can drop substantially during collapse, enhancing the capture rate. For a black hole of mass $M_{\rm BH}$, the Bondi rate is
@@
Here $\lambda$ is the usual dimensionless Bondi factor (order unity, depending on the equation of state), and $\eta$ is the radiative efficiency used in defining $\dot{M}_{\rm Edd}$. The ratio to the Eddington rate is
@@
indicating that inflow at the capture scale can exceed $\dot{M}_{\rm Edd}$ in dense cores. In the soliton scenario, the compact potential well both confines baryons within $r\sim r_{c}$ (raising $n$ for a given loaded mass) and sets the dynamical boundary condition for collapse, so that once cooling reduces $c_{s}$ the Bondi rate enhancement in Eq. (10) is naturally realized at the LRD scale; in a more weakly bound BH-only or extended-halo configuration the same column would correspond to a less compact reservoir and a smaller capture enhancement. Granule-driven fluctuations (Quantum Bondi Boost) may further modulate this (Chiu et al., 2025).
Figure 2: Timescale hierarchy in soliton cores. Comparison between the dynamical time $t_{\rm dyn}$ (blue solid line, $t_{\rm dyn}\propto M_{s}^{-2}$) and the optically thin cooling time $t_{\rm cool}$ (red dashed line; this corresponds strictly to the optically thin cooling time defined in Equation (5), not the effective diffusion-limited time $t_{\rm rad}$) as a function of soliton mass $M_{s}$. The top axis explicitly provides the corresponding virial temperature $T_{\rm vir}$ to guide physical intuition. At low masses ($T_{\rm vir}<10^{4}\,{\rm K}$), the exponential suppression of atomic cooling in primordial gas leads to a stable regime where $t_{\rm cool}>t_{\rm dyn}$. As the soliton mass increases, the gas is heated above the atomic cooling threshold, causing a sharp drop in $t_{\rm cool}$. A critical mass emerges at $M_{\rm crit}\approx 2.8\times 10^{7}M_{\odot}$, identified exactly at the intersection where the condition $t_{\rm cool}=t_{\rm dyn}$ is satisfied; this threshold is obtained analytically by solving $t_{\rm cool}=t_{\rm dyn}$ using the cooling function adopted in our model (see Appendix B). Above this mass threshold, the system enters the cooling-dominated regime ($t_{\rm cool}<t_{\rm dyn}$). For the massive regime relevant to Little Red Dots ($M_{s}\gtrsim 10^{8}M_{\odot}$), the cooling time is orders of magnitude shorter than the dynamical time, inevitably triggering a catastrophic collapse.
II.4 Thermodynamic Instability and the Cooling Catastrophe
Figure 2 summarizes the hierarchy between the dynamical timescale $t_{\rm dyn}$ and the radiative cooling timescale $t_{\rm cool}$ within the soliton core. Here we provide a scaling argument for why a crossover $t_{\rm cool}<t_{\rm dyn}$ is unavoidable at sufficiently large soliton masses.
Potential-defined dynamical time.
The relevant dynamical timescale at radius $r$ is set by the enclosed gravitating mass,
$t_{\rm dyn}(r)\equiv\sqrt{\frac{r^{3}}{G\,M_{\rm enc}(r)}}.$ (11)
Evaluated at the core scale ($r\sim r_{c}$) in the soliton-dominated regime, we estimate
$t_{\rm dyn}(r_{c})\sim\sqrt{\frac{r_{c}^{3}}{G\,M_{\rm enc}(r_{c})}}\;\simeq\;\sqrt{\frac{r_{c}^{3}}{G\,[M_{\rm BH}+{\cal O}(1)M_{s}]}}.$ (12)
Using the soliton scaling $r_{c}\propto m_{22}^{-2}M_{s}^{-1}$ (Schive et al., 2014a), for fixed particle mass $m_{22}$ this implies $t_{\rm dyn}\propto M_{s}^{-2}$ (and $t_{\rm dyn}\propto m_{22}^{-3}$ at fixed $M_{s}$).
Cooling time for a hydrostatic atmosphere.
If a gas mass $M_{g}=f_{g}f_{b}M_{s}$ is distributed within $r_{\rm e}=\xi r_{c}$, the characteristic number density scales as
$n\propto\frac{M_{g}}{r_{\rm e}^{3}}\propto\frac{f_{g}M_{s}}{(\xi r_{c})^{3}}\propto f_{g}\,\xi^{-3}\,M_{s}^{4},$ (13)
and the characteristic temperature for a hydrostatic atmosphere is set by the potential depth,
$T\sim T_{\rm vir}\propto\frac{M_{\rm enc}}{r_{c}}\propto M_{s}^{2}\quad\text{(soliton-dominated regime)}.$ (14)
The cooling time is $t_{\rm cool}\sim(nk_{B}T)/(n^{2}\Lambda)$. In the high-temperature regime where thermal bremsstrahlung is a good scaling proxy ($\Lambda\propto T^{1/2}$), this gives
$t_{\rm cool}\propto\frac{T^{1/2}}{n}\propto M_{s}^{-3}.$ (15)
Figure 2 uses the full cooling function for the numerical estimate; the scalings above are provided for intuition.
Inevitable crossover and the role of the soliton.
In primordial gas lacking molecular or metal coolants, radiative cooling is exponentially suppressed below $T\sim 10^{4}$ K due to the Boltzmann factor associated with hydrogen collisional excitation (Lyman-$\alpha$ cooling). Consequently, low-mass solitons ($M_{s}\lesssim 10^{7}M_{\odot}$, $T_{\rm vir}<10^{4}$ K) can maintain a stable, quasi-static atmosphere where the cooling time strictly exceeds the dynamical time ($t_{\rm cool}>t_{\rm dyn}$). However, as the soliton mass increases, the deeper potential well heats the gas above the $10^{4}$ K threshold, triggering highly efficient atomic cooling. In the high-temperature regime, $t_{\rm cool}$ decreases more steeply with soliton mass (e.g., $\propto M_{s}^{-3}$ for bremsstrahlung) than $t_{\rm dyn}$ ($\propto M_{s}^{-2}$). This guarantees a crossover where $t_{\rm cool}<t_{\rm dyn}$. As illustrated in the updated Figure 2, this transition defines a critical mass scale at $M_{\rm crit}\approx 2.8\times 10^{7}M_{\odot}$. The massive FDM cores relevant to LRDs ($M_{s}\gtrsim 10^{8}M_{\odot}$) already reside deeply within this cooling-dominated region. In this unstable regime, the gas loses thermal support faster than the potential-defined dynamical time, triggering rapid inflow or collapse rather than maintaining a long-lived hot atmosphere. For comparison, at the same physical radius $r\sim r_{c}$, a BH-only (“naked”) potential has $M_{\rm enc}\approx M_{\rm BH}$ and therefore a longer dynamical time than the soliton case (because $M_{\rm enc,sol}\geq M_{\rm BH}$). Thus, if $t_{\rm cool}\lesssim t_{\rm dyn}$ holds in the soliton container, it holds a fortiori in the BH-only limit. The role of the soliton is therefore not to stabilize the gas, but to define the characteristic $\sim 50$100100 pc scale relevant to LRDs and to motivate considering a quasi-spherical, volume-filling atmosphere at that scale.